# A Pioneering Study and An Innovative Information Theory-based Approach to Enhance The Transparency in Phishing Detection

### v1

@@ Results
The results in Table 3 show that our AI2TALE  method obtains the best performance on the SAC measure compared to the baselines. In particular, our AI2TALE  method achieves higher performances than the baselines from nearly 0.5% to 6.5%. These results prove the effectiveness of our method in successfully selecting the more popular and important cognitive triggers used in phishing email data through the top-$1$ selected sentences from phishing emails.
In conclusion for RQ2: The experimental results in Table 3 show the advanced capability of our AI2TALE  method in successfully learning and figuring out the commonly used and important cognitive features existing in phishing emails. That helps provide users with crucial information for understanding to defend against phishing attacks.
RQ3: Can our proposed information bottleneck theory training term and data-distribution mechanism be successful in improving the data representation learning and supporting the process of learning and figuring out the true phishing-relevant sentences in phishing emails?
@@
Table 4. The performance of our method to itself in the cases of using and without using our information bottleneck theory additional training term and the data-distribution mechanism for the Label-Accuracy and Cognitive-True-Positive measures on the testing set. We denote the Cognitive-True-Positive as Cognitive-TP for short.
Methods Label-Accuracy Cognitive-TP SAC
AI2TALE-w/oIBT2D 98.92% 98.40% 88.46%
@@
The results shown in Table 4 demonstrate the effectiveness of our proposed information bottleneck theory additional training and the data-distribution mechanism in boosting the data representation learning that facilitates and supports the process of learning and figuring out the important and true phishing-relevant information (i.e., sentences) in phishing-relevant emails.
In particular, our method consists of our introduced bottleneck theory additional training term and the data-distribution mechanism (AI2TALE) obtained better performance (from around 0.5% to 1.0%) than the case without using these terms (AI2TALE-w/oIBT2D) on Label-Accuracy, Cognitive-True-Positive, and SAC measures. In the realm of explaining machine learning when dealing with complex datasets, such as email phishing datasets, under weakly supervised conditions, these improvements are promising. That exhibits the positive effect of the bottleneck theory additional training and data-distribution mechanism terms in improving the overall performance of our AI2TALE  method to obtain state-of-the-art performance on the Label-Accuracy, Cognitive-True-Positive, and SAC measures compared to the baselines.
In conclusion for RQ3: The experimental results in Table 4 show the effectiveness of our bottleneck theory training term and the data-distribution mechanism in boosting the data representation learning process that is beneficial for not only the emails’ vulnerability prediction but also for automatically learning and figuring out the truly crucial phishing-relevant information in phishing emails compared to the case without using these terms.
@@
The training time of our framework in the cases of using and without using our proposed information bottleneck theory training term and data-distribution mechanism is considerably short, costing approximately 36.47 seconds and 21.99 seconds, respectively, for each epoch (as mentioned in Section 4.1.5, we ran our experiments on a 13th Gen Intel(R) Core(TM) i9-13900KF having 24 CPU Cores at 3.00 GHz with 32GB RAM, integrated Gigabyte RTX 4090 Gaming OC 24GB).
@@ 4.3. Human evaluation
We additionally conduct a human evaluation to investigate the usefulness of our proposed AI2TALE  method in figuring out important and phishing-related information (e.g., sentences) in phishing emails to help provide comprehensive explanations about the phishing of the corresponding email data. In particular, we evaluate whether the top-1 selected sentence in each phishing email by our proposed AI2TALE  method is perceived as convincing information for email users. To do that, we asked participants to evaluate the selected sentences of $10$ different phishing emails (randomly chosen from the testing set) in terms of whether the sentence selected in each email is important to influence and persuade users to follow the instructions in the email.
@@
4.4. Additional quantitative results
In our study, as presented in RQ1, via the Label-Accuracy metric, we aim to measure whether the selected sentences obtained from each model can accurately predict the true vulnerability label (i.e., phishing or benign) of the associated email. The intuition is that the most important sentences contribute the most to the email’s vulnerability, especially to phishing emails. In our experiments, we assess each model’s top-1 selected sentence and measure if it can effectively predict the email’s vulnerability without considering all sentences in the email. The higher the value of the Label-Accuracy measure, the better the model’s performance in selecting crucial label-relevant information from the data.
The datasets used in our paper are balanced, with an equal number of phishing and benign emails (refer to the Studied datasets section 4.1.1 for details). In the context of a balanced dataset, the accuracy measure is sufficient, and high accuracy implies low false negative (FN) and false positive (FP). To fulfill our experiments on this aspect, we calculated the F1-score for our AI2TALE  method and baselines. Specifically, we computed the F1-score in two scenarios: (i) exclusively on the testing phishing emails, and (ii) on both the testing phishing and benign emails. The results for our AI2TALE  method and the baselines (AIM, VIBI, L2X, ICVH, INVASE) in scenarios (i) and (ii) are shown in Table 5.
Table 5. The performance of our AI2TALE  method and the baselines for the F1-score measure on the testing set in two scenarios (i) and (ii).
Methods F1-score (i) F1-score (ii)
INVASE (Yoon et al., 2019a) 98.313% 98.299%
ICVH (Nguyen et al., 2021a) 96.732% 96.725%
L2X (Chen et al., 2018b) 98.261% 98.249%
VIBI (Bang et al., 2021b) 96.626%96.649%
AIM (Vo et al., 2023) 98.406% 98.399%
AI2TALE  (Ours) 99.324% 99.325%
In summary, the F1-score results for our AI2TALE  method and baselines closely align with those obtained using the accuracy measure. Notably, our AI2TALE  method exhibits consistent effectiveness and advancement, surpassing the baselines by approximately 1% to 3% in both scenarios (i) and (ii).
@@ 5. Threats to validity
Key construct validity threats are if our assessments of the proposed method and baselines demonstrate the ability for phishing attack localization. In our paper, we study an important problem of phishing attack localization where we not only aim to automatically learn and figure out the important phishing-relevant information (e.g., sentences) in each phishing email but also can detect the vulnerability $Y$ (i.e., phishing or benign) of the corresponding email based on the crucial selected information. The selected phishing-relevant information (e.g., sentences) helps provide comprehensive explanations about the vulnerability of the phishing email data. To evaluate the performance of our method and baselines, we introduce three main measures including Label-Accuracy, Cognitive-True-Positive, and SAC.
@@
Key external validity threats include whether our proposed AI2TALE  method can generalize well to different phishing vulnerabilities and other different types of phishing email datasets. We mitigated this problem by conducting our experiments on seven phishing-relevant email datasets including IWSPA-AP, Nazario Phishing Corpus, Miller Smiles Phishing Email, Phish Bowl Cornell University, Fraud emails, Cambridge, and Enron Emails.
@@ 6. Future work
While our study marks a pioneering effort and significantly contributes to advancing transparency and detection capabilities for phishing email attacks, we acknowledge the existence of one challenging case that poses a challenge to our AI2TALE  approach and other phishing detection and localization methods.
In particular, in the broad landscape of email phishing attacks, an exceptionally challenging scenario arises when phishing emails mimic the content of benign emails, differing only in the embedded URLs. For instance, attackers may send phishing emails to victims using identical text content as genuine communications from entities like PayPal, with the sole modification being the replaced URL. In this case, classifying and explaining phishing attacks primarily relying on email content and text within embedded URLs may become particularly challenging. To the best of our knowledge, this problem may need a more comprehensive approach incorporating URL analysis and additional relevant features. This augmentation aims to enhance the overall effectiveness of phishing localization models, ensuring more comprehensive and accurate detection and explanation of phishing attacks. Addressing this issue will be a focal point of our future work.
@@ 8. Conclusion
In this paper, we have successfully proposed one of the very first innovative deep learning-based methods derived from an information-theoretic perspective and information bottleneck theory for the phishing attack localization problem where automated AI-based techniques have not yet been well studied. Our AI2TALE  method can work effectively in the weakly supervised setting providing an important practical solution to solve the problem in terms of not only accurately predicting the vulnerability of the email data but also automatically learning and figuring out the important and phishing-relevant information in each phishing email. The selected information provides useful explanations about the vulnerability of the corresponding phishing email data. In addition, we also introduce appropriate measures for phishing attack localization. The rigorous and comprehensive experiments on seven real-world email datasets show the superiority of our proposed AI2TALE  method over the state-of-the-art baselines.

### v2

@@ Results
The results in Table 3 show that our AI2TALE  method obtains the best performance on the SAC measure compared to the baselines. In particular, through the most important (top-1) selected sentences from corresponding emails, our AI2TALE  method achieves higher performances than the baselines up to 6.5% (when compared to the VIBI method). These results prove the effectiveness of our method not only in figuring out the most meaningful and crucial information causing the vulnerability of the email data (as shown in Table 1) but also in consisting of the more popular and important cognitive triggers used in phishing emails through the corresponding top-$1$ selected sentences compared to the baselines.
In conclusion for RQ2: The experimental results in Table 3 demonstrate that our AI2TALE  method primarily consists of the commonly used and important cognitive principles in phishing emails compared to those of the baselines via the most important (top-1) selected sentences. That facilitates providing users with the most crucial and useful information for understanding to defend against phishing attacks.
RQ3: Can our proposed information bottleneck theory training term and data-distribution mechanism be successful in improving the data representation learning and supporting the process of learning and figuring out the most important and phishing-relevant sentences in phishing emails?
@@
Table 4. The performance of our method to itself in the cases of using and without using our information bottleneck theory additional training term and the data-distribution mechanism for the Label-Accuracy (Label-Acc) and Cognitive-True-Positive (Cognitive-TP) measures on the testing set (the higher results in bold).
Methods Label-Acc Cognitive-TP SAC
AI2TALE-w/oIBT2D 98.92% 98.10% 88.61%
@@
The results in Table 4 demonstrate the effectiveness of our information bottleneck theory additional training term and the data-distribution mechanism in boosting the data representation learning that facilitates and supports the process of learning and figuring out the most important and true phishing-relevant information (i.e., sentences) in the corresponding emails.
Although working within the context of only utilizing the top-1 selected sentences from thousands of diverse emails, particularly under the weakly supervised setting, these terms can still help improve the performance of AI2TALE  in all measures including Label-Accuracy, Cognitive-True-Positive, and SAC. The AI2TALE  method (with our information bottleneck theory additional training term and the data-distribution mechanism) approaches 99.5% for Label-Accuracy and achieves around 99% for Cognitive-True-Positive. These improvements are valuable. The results help contribute to the overall performance of our AI2TALE  method to obtain state-of-the-art performance on the Label-Accuracy, Cognitive-True-Positive, and SAC measures compared to the baselines as shown in Table 1.
In conclusion for RQ3: The experimental results in Table 4 show the effectiveness of our bottleneck theory training term and the data-distribution mechanism in boosting the data representation learning process that is beneficial for not only the emails’ vulnerability prediction but also for automatically figuring out the truly crucial and phishing-relevant information (via the most important one) in phishing emails compared to the case without using these terms.
@@
The training time of our model in the cases of using and without using our proposed information bottleneck theory training term and data-distribution mechanism is considerably short, costing approximately 36.47 seconds and 21.99 seconds, respectively, for each epoch (as mentioned in Section 4.1.5, we ran our experiments on a 13th Gen Intel(R) Core(TM) i9-13900KF having 24 CPU Cores at 3.00 GHz with 32GB RAM, integrated Gigabyte RTX 4090 Gaming OC 24GB).
@@ 4.3. Human evaluation
We additionally conduct a human evaluation to investigate the usefulness of our proposed AI2TALE  method in figuring out important and phishing-related information (e.g., sentences) in phishing emails to help provide comprehensive explanations about the phishing of the corresponding email data. In particular, we evaluate whether the most important (top-1) selected sentence in each phishing email by our proposed AI2TALE  method is perceived as convincing information for email users. To do that, we asked participants to evaluate the selected sentences of $10$ different phishing emails (randomly chosen from the testing set) in terms of whether the sentence selected in each email is important to influence and persuade users to follow the instructions in the email.
@@ 5. Threats to validity
Key construct validity threats are if the assessments of our proposed method and baselines demonstrate the ability for phishing attack localization. In our paper, we study an important problem of phishing attack localization where we not only aim to automatically learn and figure out the most important and phishing-relevant information (e.g., sentences) in each phishing email but also can detect the vulnerability $Y$ (i.e., phishing or benign) of the corresponding email based on the crucial selected information. The selected phishing-relevant information (e.g., sentences) helps provide useful and concise explanations about the vulnerability of the phishing email data. To evaluate the performance of our method and baselines, we use three main measures including Label-Accuracy, Cognitive-True-Positive, and SAC.
@@
Key external validity threats include whether our proposed AI2TALE  method can generalize well to different types of phishing email datasets. We mitigated this problem by conducting our experiments on seven real-world diverse phishing-relevant email datasets including IWSPA-AP, Nazario Phishing Corpus, Miller Smiles Phishing Email, Phish Bowl Cornell University, Fraud emails, Cambridge, and Enron Emails.
@@ 6. Future work
In this study, our paper focuses primarily on addressing the phishing attack localization problem in email phishing, the most common form of phishing. It is worth noting that phishing attacks can also occur through webpages. Several methods have been proposed to detect and infer phishing intention based on webpage appearances (e.g., (Abdelnabi et al., 2020b; Liu et al., 2022)). We believe the principles underlying our AI2TALE  framework can also be extended to detect and explain webpage phishing attacks. The operational nature of our AI2TALE approach within (i) a unified framework, (ii) working directly with webpage data without requiring additional steps to gain extra information (e.g., references to the ground truth of phishing information) beyond the data and its vulnerability label (i.e., phishing or benign), as well as (iii) the way our AI2TALE model can be trained without requiring the ground truth of phishing information in the data can be considered as some of the advantages of our framework compared to the relevant methods (e.g., (Abdelnabi et al., 2020b; Liu et al., 2022)). Investigating the application of our AI2TALE  framework to explain webpage phishing attacks could be a focus of our future studies.
@@ 8. Conclusion
In this paper, we have successfully proposed an innovative deep learning-based method derived from an information-theoretic perspective and information bottleneck theory for solving the phishing attack localization problem where automated AI-based techniques have not yet been well studied. Our AI2TALE  method can work effectively in the weakly supervised setting providing an important practical solution to solve the problem in terms of not only accurately predicting the vulnerability of the email data but also automatically learning and figuring out the most important and phishing-relevant information in each phishing email. The selected information provides useful and concise explanations about the vulnerability of the corresponding phishing email data. In addition, we also introduce appropriate measures for phishing attack localization. The rigorous and comprehensive experiments on seven real-world diverse email datasets show the superiority of our proposed AI2TALE  method over the state-of-the-art baselines.